# Oho 1-1, Tsukuba, Ibaraki 305-0801, Japan

### v1

@@ 5 Summary and Conclusion
This paper introduces an analysis model (AM) employing the important high-level features (HLFs) in top jet tagging. The AM simplifies the current state-of-the-art taggers analyzing low-level features (LLFs) while maintaining classification performance. To demonstrate the effectiveness of our model, we revisited the problem of top jet tagging using various Monte Carlo simulated samples at the angular resolution of a hadronic calorimeter scale. We compared the classification performance of AM with a classifier based on the Particle Transformer (ParT). In our previous works, we have developed AMs using the IRC-safe relation network Lim:2018toa ; Chakraborty:2019imr ; Chakraborty:2020yfc and mathematical morphology for jet physics Chakraborty:2020yfc ; Lim:2020igi . This model’s performance matched the performance of a convolutional neural network, but it fell short of matching ParT’s performance. Nevertheless, once we fully considered the newly introduced features – $p_{T}$/̄conditioned jet constituent multiplicity distribution and subjet constituent multiplicities for capturing color charges of subjets – the performance of AM is aligned with that of ParT within a range of $2\sim 3\sigma$ of training uncertainties.
@@ Acknowledgements.
We thank David Shih for useful discussions. This work is supported by Grant-in-Aid for Transformative Research Area (A) 22H05113 and Grant-in-Aid for Scientific Research(C) JSPS KAKENHI Grant Number 22K03629. The work of SHL was also supported by the US Department of Energy under grant DE-SC0010008. The authors acknowledge the Office of Advanced Research Computing (OARC) at Rutgers, The State University of New Jersey for providing access to the Amarel cluster and associated research computing resources that have contributed to the results reported here (URL: https://oarc.rutgers.edu). This paper is revised using a large language model, ChatGPT4. The authors guarantee that the AI tool is used only to improve writing quality, not to generate the idea and paper itself.
@@ Appendix A Classification Results Involving Vincia Simulations
Appendix B Comparing Computational Resources utilized by AM and ParT
In this appendix, we describe the computation resources utilized by AM and ParT. During training with a batch size of 1,000, AM requires less than 1 GB of GPU memory, while ParT consumes approximately 14 GB. Additionally, the training process of AM is significantly faster. When training AM on an NVIDIA GeForce 1080 Ti GPU (FP32 performance: 11.3 TFLOPS), the training session takes 30 minute with an overall GPU utilization of 35%. In comparsion, training of ParT on an NVIDIA RTX 6000 GPU (FP32 performance: 38.7 TFLOPS) also takes 30 minutes, but with a GPU utilization of 95%. The ratio of the number of floating-point operations, calculated as (FP32 performance) $\times$ (GPU utilization) $\times$ (training time), between ParT and AM is approximately 18.3 to 2.0. This indicates that AM requires a significantly smaller number of GPU operations than ParT.
Appendix C Proof of Classifier Interpretation of Odds Ratio

### v2

@@ 6 Summary and Conclusion
This paper introduces an analysis model (AM) employing the important high-level features (HLFs) in top jet tagging. The AM simplifies the current state-of-the-art taggers analyzing low-level features (LLFs) while maintaining classification performance. To demonstrate the effectiveness of our model, we revisited the problem of top jet tagging using various Monte Carlo simulated samples at the angular resolution of a hadronic calorimeter scale. We compared the classification performance of AM with a classifier based on the Particle Transformer (ParT). In our previous works, we have developed AMs using the IRC-safe relation network Lim:2018toa ; Chakraborty:2019imr ; Chakraborty:2020yfc and mathematical morphology for jet physics Chakraborty:2020yfc ; Lim:2020igi . This model’s performance matched the performance of a convolutional neural network, but it fell short of matching ParT’s performance. Nevertheless, once we fully considered the newly introduced features – $p_{T}$-conditioned jet constituent multiplicity distribution and subjet constituent multiplicities for capturing color charges of subjets – the performance of AM is aligned with that of ParT within a range of $2\sim 3\sigma$ of training uncertainties.
@@ Acknowledgements.
We thank David Shih and Edward Ramirez for the useful comments and discussions. This work is supported by Grant-in-Aid for Transformative Research Area (A) 22H05113 and Grant-in-Aid for Scientific Research(C) JSPS KAKENHI Grant Number 22K03629. The work of SHL was also supported by the US Department of Energy under grant DE-SC0010008. The authors acknowledge the Office of Advanced Research Computing (OARC) at Rutgers, The State University of New Jersey for providing access to the Amarel cluster and associated research computing resources that have contributed to the results reported here (URL: https://oarc.rutgers.edu). This paper is revised using large language models, ChatGPT4 and Claude 3 Opus. The authors guarantee that the AI tools are used only to improve writing quality, not to generate the idea and paper itself.
@@ Appendix A Classification Results Involving Vincia Simulations
Appendix B Maximum Dilation Scale Choice in Morphology Analysis Module
The morphology analysis module in Section 2.4 has a hyperparameter: maximum dilation scale $R_{\mathrm{max}}$. In this appendix, we display the influence on classifier performance under the change of this scale parameter.
The choice of this parameter may vary the performance not only in jet tagging but also in simulation comparison since the angular distribution of parton shower pattern depends on its modeling. The subsequent emissions in parton showers with angular ordering of HW and $p_{T}$-ordering of PY are different Webber:2010vz ; Bhattacherjee:2015psa , and the influence can also be observed by two-point energy correlation $S_{2}$ based analysis Chakraborty:2019imr . On top of this, the Minkowski functionals capture the geometric features of parton shower in angular coordinate $(\eta,\phi)$, which are more sensitive to soft physics compared to the energy correlations.
Figure 8: Left: The $R_{\mathrm{max}}$ dependence of $R_{30\%}$ of the $\mathrm{P_{T}}$ vs. $\mathrm{H_{T}}$ classification (blue lines) and $\mathrm{P_{Q}}$ and $\mathrm{H_{Q}}$ classification (red lines). The solid lines are for the AM, and the dashed lines are for the AM-PIP E) $x_{\mathrm{kin}}$, $x_{S_{2}}$, $x_{\mathrm{MF}}$ + $x_{\mathrm{count}}$. The $R_{30\%}$ of the ParT model is also shown in horizontal lines. All errors are estimated by bootstrapping. Right: The same plot for $\mathrm{P_{T}}(\mathrm{H_{T}})$ vs. $\mathrm{P_{Q}}(\mathrm{H_{Q}})$ classification.
In Figure 8, we show the rejection rates $R_{30\%}$ of the AM for different choices of $R_{\mathrm{max}}$. The left plots show the PY vs. HW classification result for top and QCD jets. Each point in the plot is the average and standard deviation of $R_{30\%}$ for the different bootstrap datasets. We can see that too small $R_{\mathrm{max}}$ results in degraded performance since topological and geometrical features prominent at a larger dilation scale are not counted yet, but as $R_{\mathrm{max}}$ increases, the rejection rate approaches that of ParT and starts to saturate from $R_{\mathrm{max}}=0.3$
In the same plot, we also show in dashed line the $R_{\mathrm{max}}$ dependence of AM-PIP E) taking the following inputs: $x_{\mathrm{kin}}$, $x_{S_{2}}$, $x_{\mathrm{count}}$ and $x_{\mathrm{MF}}$. AM shows a weaker dependence on $R_{\mathrm{max}}$ compared to AM-PIP E). This is because $x_{\mathrm{MF}}$ and the subjet constituent multiplicies in $x_{\mathrm{subj}_{ex}}$ for $R=0.1,0.2,$ and $0.3$ are partially correlated. The residual $R_{\mathrm{max}}$ dependence of $R_{30\%}$ of AM comes from the jet constituents outside the leading subjets. The variances of rejection rates of AM under the scaling of $R_{\mathrm{max}}$ is around 5% for AM and 10% AM-PIP E) within the range [0.025, 0.3], and they are also statistically significant.
As discussed in the main text, the estimated error of $R_{30\%}$ for the ParT model is significantly larger than that for AM(-PIP) models. The solid and dashed horizontal lines show $R_{30\%}$ of ParT and its $1\sigma$ deviation estimated by bootstrapping. The deviation is almost as large as the changes of $R_{30\%}$ of AM-PIP from $R_{\mathrm{max}}=0.025$ to 0.3. This illustrates that AM is beneficial in interpreting the classifiers, such as precision analysis on the influence of input features on the classification performance.
In the right plot, we show the results of $\mathrm{P_{T}}$ vs. $\mathrm{P_{Q}}$ and $\mathrm{H_{T}}$ vs. $\mathrm{H_{Q}}$ classification. Unlike the top jet tagging problems, $\mathrm{P_{Q}}$ vs. $\mathrm{H_{Q}}$ and $\mathrm{P_{T}}$ vs. $\mathrm{H_{T}}$, we do not see significant $R_{\mathrm{rmax}}$ dependences. This is expected because the difference of the hard substructure is the dominant part of the classification, and the parton shower emissions and the hadronization model are common between $\mathrm{P_{T}}$ and $\mathrm{P_{Q}}$ or $\mathrm{H_{T}}$ vs. $\mathrm{H_{Q}}$.
Appendix C Comparing Computational Resources utilized by AM and ParT
In this appendix, we describe the computation resources utilized by AM and ParT. During training with a batch size of 1,000, AM requires less than 1 GB of GPU memory, while ParT consumes approximately 14 GB. Additionally, the training process of AM is significantly faster. When training AM on an NVIDIA GeForce 1080 Ti GPU (FP32 performance: 11.3 TFLOPS), the training session takes 30 minutes with an overall GPU utilization of 35%. In comparison, training of ParT on an NVIDIA RTX 6000 GPU (FP32 performance: 38.7 TFLOPS) also takes 30 minutes, but with a GPU utilization of 95%. The ratio of the number of floating-point operations, calculated as (FP32 performance) $\times$ (GPU utilization) $\times$ (training time), between ParT and AM is approximately 18.3 to 2.0. This indicates that AM requires a significantly smaller number of GPU operations than ParT.
Appendix D Proof of Classifier Interpretation of Odds Ratio